# MoCA: Multi-modal Cross-masked Autoencoder for Digital Health Measurements

### v1

@@ 4.4 Connection to Canonical Correlation Analysis (CCA)
4.5 Influence of Masking Schemes
Since any random masking scheme defines a probability measure over the space of admissible masking realizations of $\delta$, the objective in (5) can be interpreted as a weighted average of the optimal values of the objective in (6). If the matrices $K_{u}$ and $K_{m}$ are known or can be reasonably approximated, then the optimal random masking scheme is one that assigns higher probabilities to masking allocations that yield larger objective values.
In the special case of linear kernels, the optimal objective is the top singular value of the matrix $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ (Sánchez et al., 1979; Andrew et al., 2013). Therefore, to optimize (5), higher probabilities should be assigned, subject to the mask ratio constraint, to masking allocations that result in matrices $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ with larger top singular values.
The matrix $\Sigma^{-1/2}_{UU}\Sigma_{UM}\Sigma^{-1/2}_{MM}$ can be interpreted as a normalized adjacency matrix of a bipartite graph, where nodes are masked and unmasked patches, and edge weights are values in the corresponding elements of the matrix. A large top singular value indicates a potentially strong community/cluster of nodes that are highly correlated with each other (Sarkar & Dong, 2011), that is, pairs of masked and unmasked patches with large absolute values of the correlation. Intuitively, to optimize (5), the random masking scheme should generate such pairs with larger probabilities.
4.6 Demonstration with Synthetic Data
To validate the above heuristic, we compare MoCA against synchronized masking through simulations with synthetic data. Following Algorithm 1, we generate synthetic data by sampling two samples $\bar{x}_{1},\bar{x}_{2}$ from the training set, and then randomly extract a consecutive chunk of time series from $\bar{x}_{2}$ to replace the corresponding time series in $\bar{x}_{1}$. We repeated this process 1000 times to obtain our simulation data. An example image is shown in Figure 1. With both MoCA and synchronized masking, we calculate the averaged optimal objective in (6). MoCA yields a value of 0.1989, comparing to 0.1812 with synchronized masking. This result demonstrates that MoCA generates masked and unmasked patches that better capture cross-modal correlations, leading to improved reconstruction, whereas synchronized masking fails to leverage correlations between patches from different modalities within the same time window.
@@ 5.4 Generative Task
Wearable devices have transformed health monitoring and other domains, but missing data remains a persistent challenge that limits their reliability. Data gaps can arise from signal glitches, battery depletion, or device malfunction. These challenges are magnified when multiple sensors are used. In this section, we evaluate each methods with respect to real world challenges in health monitoring and assess each model’s ability to reconstruct missing data effectively. The tasks include the following: (i) Random Imputation tests the model’s ability to reconstruct localized missing values by randomly masking small data patches and predicting their original values, (ii) Temporal Imputation evaluates long-term dependency modeling by masking all sensor data over consecutive time periods and requiring the model to predict the missing segments, and (iii) Sensor Imputation measures cross-modality learning by masking an entire sensor channel and having the model reconstruct it using information from correlated signals.
Similar to (Narayanswamy et al., 2024), we will evaluate the reconstruction results on the UCI-HAR testing set against two widely used statistical imputation methods: (1) Linear Interpolation, which fills missing values by linearly interpolating between known values along the temporal dimension, and (2) Nearest Neighbor Imputation, which imputes missing values by averaging the values of the k-nearest neighbors to handle missing data. To evaluate the imputation methods, we use Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better performance.
@@ 6 Conclusion
However, self-supervised learning can be implemented by anyone who has no access to the true label of the data, and this might pose threats to the safety of the deployed model. Once our model is published and made public, such self-supervised model without proper fintuning or professional interpretations may possibly be misleading in its prediction and implications. We recommend careful and thoughtful application and reproduction of our methods.

### v2

@@ 5.4 Generative Task
Wearable devices have transformed health monitoring and other domains, but missing data remains a persistent challenge that limits their reliability. Data gaps can arise from signal glitches, battery depletion, or device malfunction. These challenges are magnified when multiple sensors are used. In this section, we evaluate each methods with respect to real world challenges in health monitoring and assess each model’s ability to reconstruct missing data effectively. The tasks include the following: (i) Random Imputation tests the model’s ability to reconstruct localized missing values by randomly masking small data patches and predicting their original values, (ii) Temporal Imputation evaluates long-term dependency modeling by masking all sensor data over consecutive time periods and requiring the model to predict the missing segments, and (iii) Sensor Imputation measures cross-modality learning by masking an entire sensor channel and having the model reconstruct it using information from correlated signals, (iv) Temporal Extrapolation evaluates the model’s ability to forecast future values by masking all sensor data beyond a certain time point and requiring the model to predict the upcoming sequence based only on past observations.
Similar to (Narayanswamy et al., 2024), we will evaluate the reconstruction results on the UCI-HAR testing set against two widely used statistical imputation methods: (1) Linear Interpolation, which fills missing values by linearly interpolating between known values along the temporal dimension, and (2) Nearest Neighbor Imputation, which imputes missing values by averaging the values of the k-nearest neighbors to handle missing data. We additionally experimented with Multivariate Imputation by Chained Equations (MICE) (Buuren & Groothuis-Oudshoorn, 2011) as a supplementary baseline. However, both MICE and nearest neighbor imputation struggle under large masking ratios, leading to highly similar outputs. To maintain visual clarity and avoid redundancy, we exclude MICE from the main visualizations. To evaluate the imputation methods, we use Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better performance.
@@ 6 Conclusion
However, self-supervised learning can be implemented by anyone who has no access to the true label of the data, and this might pose threats to the safety of the deployed model. Once our model is published and made public, such self-supervised model without proper finetuning or professional interpretations may possibly be misleading in its prediction and implications. We recommend careful and thoughtful application and reproduction of our methods.
7 Acknowledgements
This project was funded by NIH/NHLBI grant R01HL166802.